OpenStudy (prince1342):

1. The main tank has a radius of 70 feet. What is the volume of the quarter-sphere sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (prince1342):

@563blackghost

563blackghost (563blackghost):

So to find the volume of a sphere we use this equation.... \(\Huge{V=\frac{4}{3} \pi r^{3}}\) Where \(\large{r}\) is \(\large{70}\)...so input.. \(\Huge\color{purple}{V=\frac{4}{3} (3.14) 70^{3}}\) Simplify....

563blackghost (563blackghost):

Once we find the volume of a whole sphere we would then divide it by four to get the volume of a quarter-sphere tank....

OpenStudy (prince1342):

Hmmmm I think I get it

563blackghost (563blackghost):

So you were able to come up with your answer?

OpenStudy (prince1342):

plugging it all in now, sorry if its taking me a while my computer is so slow today

563blackghost (563blackghost):

its alright ^^

OpenStudy (prince1342):

I got 0.0000122060252672497 as my answer lol

563blackghost (563blackghost):

\(\Huge{V=\frac{4}{3}(1,077,020)}\) We multiply it by 4.... \(\Huge\color{orange}{V=\frac{4,308,080}{3}}\) In which you would divide by 3....

OpenStudy (prince1342):

1436026.6666666666666666666666667

OpenStudy (prince1342):

Is this right?

563blackghost (563blackghost):

Yup ^^ Now we need to find a quarter of a sphere so we would divide this number by 4.... \(\Huge{\frac{1436026.667}{4}}\)

OpenStudy (prince1342):

Why would we divide by 4?

OpenStudy (prince1342):

oh and I got 359006.66666666666666666666666667

563blackghost (563blackghost):

Correct but they want you to round to the nearest who so it would round to \(\Huge{359,007}\) Since 60 is closer to 100....

OpenStudy (prince1342):

Ohhhh ok, thanks! I'll explain how I came to that on my assingment, I'll post my next/last question in a bit.

563blackghost (563blackghost):

glad i was helpful ^^ just tag me in your next question :)